### v1

# MedArabiQ: Benchmarking Large Language Models on Arabic Medical Tasks

Abstract
Large Language Models (LLMs) have demonstrated significant promise for various applications in healthcare. However, their efficacy in the Arabic medical domain remains unexplored due to the lack of high-quality domain-specific datasets and benchmarks. This study introduces MedArabiQ, a novel benchmark dataset consisting of seven Arabic medical tasks, covering multiple specialties and including multiple choice questions, fill-in-the-blank, and patient-doctor question answering. We first constructed the dataset using past medical exams and publicly available datasets. We then introduced different modifications to evaluate various LLM capabilities, including bias mitigation. We conducted an extensive evaluation with five state-of-the-art open-source and proprietary LLMs, including GPT-4o, Claude 3.5-Sonnet, and Gemini 1.5. Our findings highlight the need for the creation of new high-quality benchmarks that span different languages to ensure fair deployment and scalability of LLMs in healthcare. By establishing this benchmark and releasing the dataset, we provide a foundation for future research aimed at evaluating and enhancing the multilingual capabilities of LLMs for the equitable use of generative AI in healthcare.
Data and Code Availability
In this article, we present a new benchmark dataset, MedArabiQ, and evaluate state-of-the-art LLMs. We release our data for reproducibility, to enable a fair evaluation of LLM models in the future: https://github.com/nyuad-cai/MedArabiQ
Institutional Review Board (IRB)
This study does not involve human subjects, so IRB approval was not required.
1 Introduction
The recent advent of Large Language Models (LLMs) has revolutionized Natural Language Processing (NLP) by demonstrating exceptional performance in various tasks, ranging from language translation to creative writing (Nazi and Peng, 2023). Although LLMs were initially designed for general language understanding, they have since been evaluated for many domain-specific applications, such as education, programming, art, and medicine. They have also been adapted for domain-specific tasks using different fine-tuning strategies and specialized datasets (Kaddour et al., 2023).
The use of LLMs in healthcare has sparked enthusiasm due to the potential to improve diagnostic processes, clinical decision-making, and overall patient care (Nazi and Peng, 2023; Meng et al., 2024). One particular application of interest is medical education, in which LLMs can generate concise summaries of information and support interactive learning experiences (Benítez et al., 2023). To this end, several benchmarks have been proposed to assess the capabilities of LLMs in medical knowledge and reasoning. Despite these advances, challenges remain, including ethical concerns, risks of generating biased or harmful content, and variability in performance in different languages and cultural contexts (Yang et al., 2024b; Nazi and Peng, 2023).
Existing benchmarks, such as GLUE and MedQA, cater primarily to English, leaving a significant gap in the evaluation of LLMs for Arabic healthcare applications (Qiu et al., 2024). This is due to multiple factors including the limited availability of high-quality Arabic datasets for clinical applications, coupled with the unique linguistic complexity of Arabic, especially when considering the many different Arabic dialect regions (e.g., Gulf, Maghreb, Egypt, Levant, among others) in addition to Modern Standard Arabic (Salameh et al., 2018). Additionally, although various multilingual models include Arabic in their training data, their performance often falls short in clinical contexts due to insufficient domain-specific resources and a lack of appropriate benchmarks (Nazi and Peng, 2023; Gangavarapu, 2024). Addressing these gaps is critical to unlocking the full potential of LLMs for Arabic-speaking patients and providers, ensuring equitable access to AI-driven advancements in healthcare.
To address these gaps, there is a growing need for frameworks that assess LLM performance in clinical tasks specific to Arabic-speaking populations. By developing benchmarks that reflect real-world clinical interactions, we can ensure more reliable and culturally appropriate LLM deployment in multilingual healthcare systems. In this study, we make several key contributions to address these challenges by introducing MedArabiQ (see Figure 1). First, we developed seven benchmark datasets designed to evaluate LLMs in Arabic healthcare applications while addressing Arabic linguistic complexity along with the domain-specific challenges hindering existing models. We focused on critical healthcare medical tasks, including medical question answering, clinical dialogue, and ethical decision-making. Second, we analyzed the performance of both multilingual and Arabic LLMs, highlighting the impact of linguistic coverage and training data transparency on healthcare applications. We conducted comprehensive evaluation to assess model performance, in order to provide a robust foundation for advancing AI-driven solutions for Arabic healthcare tasks.
Figure 1: Overview of MedArabiQ. We constructed seven new benchmark datasets and evaluated five state-of-the-art LLMs on them.
2 Related Work
Several benchmarks have been proposed for the evaluation of LLMs for various medical tasks, predominantly in the English language. For example, Gao et al. (2023) introduce Dr. Bench, a diagnostic reasoning benchmark for clinical NLP emphasizing clinical text understanding, medical knowledge reasoning, and diagnosis generation. The benchmark combines English-language data from various sources but primarily consists of in-hospital clinical notes. To assess LLM performance in medical question-answering tasks, a number of benchmarks incorporated samples from medical board exams, such as MedQA. MedQA advances multilingual benchmarking by including questions in traditional and simplified Chinese, and English (Jin et al., 2020). The Massive Multitask Language Understanding (MMLU) benchmark uses the US Medical Licensing Exam (MLE) (Hendrycks et al., 2021) for a subset of tasks. MedMCQA extends these benchmarks to a multilingual evaluation framework (Pal et al., 2022).
Despite the growing interest in Arabic NLP, Arabic medical benchmarks remain limited. Achiam et al. (2023) translated MMLU into 14 languages, including Arabic, with the help of professional human translators. AraSTEM focuses on the task of question-answering and has a medical subset (Mustapha et al., 2024). AraMed similarly presents an Arabic medical corpus and annotated Arabic question-answering dataset (Alasmari et al., 2024). Other datasets are also largely skewed towards the task of Arabic medical question-answering and often have other shortcomings, summarized in Table A1. While these resources are helpful, they do not comprehensively cover the spectrum of Arabic medical tasks, highlighting the need for dedicated benchmarking efforts.
In terms of evaluation, a number of frameworks have been proposed for evaluating clinical AI models. The ‘Governance Model for AI in Healthcare’ consists of four main pillars: fairness, transparency, trustworthiness, and accountability (Reddy et al., 2020). Similarly, Dada et al. (2024) propose the Clinical Language Understanding Evaluation framework, which was designed to assess LLMs with real patient data for various modalities, including patient-specific question-answering, hypothesis deduction, and problem and inquiry summarization. Kanithi et al. (2024) introduce ‘MEDIC’, a framework for evaluating LLMs across five clinically relevant categories: medical reasoning, ethical and bias concerns, data and language understanding, in-context learning, and clinical safety and risk assessment. Although there has been significant progress in the field of benchmarking LLMs in medical tasks, most of the evaluations have been conducted for data in the English language, and no single benchmark exists to assess LLMs in Arabic on more than one medical task and performance metric. There are over 380 million native Arabic speakers (Eberhard et al., 2024), a number of which are monolingual. With the vast potential of LLMs in healthcare, it is crucial to accommodate Arabic-speaking patients to ensure fair deployment.
3 Methodology
Here we describe the details of our methodological framework to construct the datasets and evaluate state-of-the-art LLMs. An overview of MedArabiQ is provided in Figure 1.
3.1 Tasks and Datasets
In order to develop a reliable framework for evaluating LLMs in Arabic healthcare applications, we focused on telehealth consultation and question-answering as the main use cases. These involve not only medical reasoning ability but also natural patient-doctor dialogue. The LLM should mimic the role of a clinician as closely as possible, which includes possession of medical knowledge and the ability to utilize the knowledge, personalized to the needs and queries of patients. Personalization, however, should not lead to bias or prejudice in the LLM’s response, based on the patient’s profile. In fact, the model should show resistance to biased prompts, responding with fairness and objectivity. We derived our datasets from two primary sources: past exams and notes from Arabic medical schools, and the AraMed Dataset (Alasmari et al., 2024). We specifically selected data sources that were unlikely to have been included in prior training datasets.
3.1.1 Multiple Choice Questions
To evaluate the models’ medical understanding, we curated a standard dataset with question-answer pairs, covering foundational and advanced medical topics, such as physiology, anatomy, and neurosurgery. First, we collected paper-based past exams and lecture notes sourced from a large repository of academic materials hosted on student-led social platforms of regional medical schools. No personally identifiable information or real patient data was included, and thus anonymization was not necessary as our data collection complied with privacy and ethical guidelines. These exams and answers were not readily available in structured digital formats, necessitating a rigorous manual process to ensure clarity and correctness. Given that Arabic medical education is not widely digitized, these exams are not publicly accessible in structured formats. Even if some individual questions exist online, the extensive effort required to compile, format, and structure them into a benchmark dataset significantly reduces the likelihood of contamination. Questions were selected to reflect increasing complexity across different academic years, ensuring that model performance could be assessed at varying levels of medical expertise. We selected a random set of 100 multiple-choice questions, digitized and extracted them into CSV files, and performed manual verification. The average question length is 15-30 words.
3.1.2 Multiple Choice Questions with Bias
Following recent work (Schmidgall et al., 2024), we injected bias in multiple choice questions dataset to evaluate how LLMs handle ethical or culturally sensitive scenarios. In particular, we utilized a set of well-defined bias categories proposed by Schmidgall et al. (2024), including (i) confirmation bias, (ii) recency bias, (iii) frequency bias, (iv) cultural bias, (v) false-consensus bias, (vi) status quo bias, and (vii) self-diagnosis bias. By manually injecting the bias, we ensured relevance to the unique linguistic and clinical challenges of Arabic healthcare contexts. This resulted in a dataset consisting of 100 samples.
3.1.3 Fill-in-the-Blank with Choices
To assess knowledge recall and in-context learning, we manually constructed fill-in-the-blank questions, each accompanied by a set of predefined answer choices. The model was required to select the most appropriate answer from the given options. This approach evaluates the model’s ability to recognize correct answers within a constrained set, reducing the reliance on generative capabilities. The resulting dataset consists of 100 samples.
3.1.4 Fill-in-the-Blank without Choices
In this setting, the fill-in-the-blank questions were presented without predefined answer choices, requiring the model to generate responses independently. This evaluation measures the model’s ability to recall and generate accurate medical knowledge without additional information, emphasizing its reasoning and language generation capabilities. The dataset for this task also comprises 100 samples.
3.1.5 Patient-doctor Q&A
AraMed is an Arabic medical corpus for question-answering that was originally sourced from AlTibbi, an online medical patient-doctor discussion forum (Alasmari et al., 2024). The original dataset consists of 400 question-answer pairs. These are all already formulated in the original corpus in Modern Standard Arabic (MSA), ensuring consistency and standardization. We manually selected 100 samples. Our selection process prioritized questions with well-formed queries and meaningful answers, avoiding instances where responses were overly generic (e.g., “Consult a doctor” or “See a specialist”). However, we retained some examples of such cases to reflect real-world user behavior, as patients often seek medical advice even for questions that are better suited for in-person consultations.
Additionally, we maintained proportionality in representation to mirror the dataset’s original structure, particularly for categories like reproductive and sexual health. These categories comprised a significant portion of the dataset and are often underrepresented in Arabic-language medical research despite their importance. Given the sensitivity and cultural taboos associated with these topics, we ensured their inclusion to provide a more realistic and balanced evaluation of the model’s ability to handle diverse medical inquiries. This approach ensures the dataset reflects the varied nature of medical concerns while maintaining its relevance to real-world healthcare scenarios. In order to ensure proportional representation, we selected 100 questions, allocating an equal proportion of samples to each specialty: cardiology, obstetrics and gynecology, surgery, pediatrics, neurology, oncology, endocrinology, dentistry, otholarhyntology, public health, dermatology, primary care, pulmonology, and psychology.
We also incorporated information about the patient, specifically age and gender, when known, into the query. When this information was unknown or erroneous (i.e., the age was an illogical number), it was excluded. Some questions came from an acquaintance of the patient, so no information was known about the patient. The information was usually prepended to the question or inserted after the greeting, in the format of “I am a [man/woman] and I am [x] years old”. The inclusion of this information was useful in incorporating personalization and realism into the case scenarios, making our benchmark dataset better suited to evaluate LLM performance in real-world medical scenarios.
3.1.6 Q&A with Grammatical Error Correction (GEC)
Since the patient Q&A dataset uses dialectal Arabic, we constructed an additional version with enhanced linguistic quality and consistency. We specifically applied a Grammatical Error Correction (GEC) pipeline tailored for Arabic healthcare texts. Given the morphological complexity and syntactic richness of Arabic, this preprocessing step was essential. First, we used cameltools, an open-source Arabic NLP library, to morphologically disambiguate words and dediacritize text, ensuring uniformity (Obeid et al., 2020). This step was crucial for handling Arabic’s inflectional patterns and preparing the dataset for grammatical correction. Then, we employed an existing fine-tuned BERT-based Arabic Grammatical Error Detection (GED) model (Alhafni et al., 2023) to detect agreement errors, incorrect word order, and missing inflections. Each token was tagged with a grammatical label to facilitate structured corrections. Next, we used the GEC model from the same pipleine (built on mBART), to automatically correct detected errors while preserving semantic integrity. The model was fine-tuned on the QALB-2015, QALB-2014, and ZAEBUC corpora to enhance its performance in grammatical corrections (Mohit et al., 2014; Rozovskaya et al., 2015; Habash and Palfreyman, 2022).
3.1.7 Q&A with LLM Modifications
Additionally, to mitigate potential model memorization, since some models could have been trained using scraped data of Altibbi, we modified the dataset using an LLM for a more rigorous assessment of reasoning and adaptability (Dong et al., 2024a). In particular, we used GPT-4o to paraphrase our original questions using the prompt, “You are a helpful assistant that paraphrases text while keeping its meaning intact.” This approach ensured that the core medical concepts remained unchanged while reducing the likelihood of models relying on memorized content.
In summary, we constructed seven new datasets using two primary sources (AraMed and past medical exams) via extensive manual verification to build the MedArabiQ benchmark.
[…]
tab:benchmark_results Benchmark Datasets Metrics LLaMA Claude GPT-4o Qwen Gemini Multiple Choice Questions Accuracy 26.2 53.5 53.5 38.0 57.5 Fill-in-the-Blank with choices Accuracy 42.4 72.7 70.7 42.40 72.7 Fill-in-the-Blank without choices BERTScore 78.2 83.6 88.2 83.5 88.8 Patient-Doctor Q&A BERTScore 81.2 81.1 84.5 85.2 82.5 Q&A with Grammatical Error Correction BERTScore 85.5 83.7 84.5 85.2 82.5 Q&A with LLM modifications BERTScore 84.9 83.6 84.2 84.9 82.3
Table 1: Summary of performance results across six benchmark datasets. We present the results in terms of accuracy and BERTScore, depending on task, and show best results in bold per row (Note we evaluate LLaMA 3.1, Claude-3.5-Sonnet, GPT-4o, Qwen 2.5, and Gemini 1.5 Pro).
3.2 Models
Benchmarks used to evaluate LLMs often face the challenge of data contamination. Many modern LLMs are trained on large web crawl datasets sourced from the internet, which can inadvertently include benchmark questions. This overlap inflates performance metrics, making such benchmarks unreliable for fair evaluation (Chunyuan et al., 2024; Dong et al., 2024b). Addressing this issue is critical to ensure that benchmark results accurately reflect the true capabilities of LLMs, free from biases introduced by prior exposure to evaluation data.
To mitigate these concerns, we categorize models into two primary groups: those with known training data and those with unknown training data. This initial distinction allows us to assess contamination risks more reliably for models whose training sources are documented, while acknowledging that for models without such transparency, contamination cannot be easily ruled out.
In addition to categorizing models based on the transparency of their training data, this study further classifies models according to their linguistic coverage. Specifically, we distinguish between models that are broadly multilingual and those that are primarily Arabic-centric multilingual. Although many large language models include training data from numerous languages, exposure to a language does not inherently confer proficiency (Bender et al., 2021). GPT-4, for example, is known to have been trained on a variety of languages (Achiam et al., 2023). By assessing multilingual and Arabic-centric multilingual models separately, we determine how language coverage and specialization influence performance, particularly in the domain of Arabic medical tasks. Table C1 shows the models selected.
3.3 Evaluation
The evaluation of LLMs in Arabic healthcare applications requires a comprehensive framework that balances technical performance with real-world applicability (Sallam and Mousa, 2024). Our framework assesses medical reasoning, decision-making, and dialogue-based question answering (Kanithi et al., 2024; Guo et al., 2023). Multiple choice questions were evaluated based on accuracy, while open-ended questions, such as those in the fill-in-the-blank and patient-doctor Q&A, were assessed using BERTScore to measure semantic alignment (Zhang* et al., 2020).
3.4 Bias Assessment and Mitigation
To ensure that LLMs can be effectively deployed in healthcare and provide meaningful contributions to the field, it is critical to address their potential to replicate human biases. Completely eliminating bias from LLMs is nearly impossible, as the datasets used to train these models are inherently shaped by human judgment, which is itself susceptible to bias. Recognizing this challenge, we developed an evaluation framework to systematically assess model resistance to biases injected into the multiple choice questions, measure susceptibility, and test mitigation strategies. The framework draws on methodologies outlined in recent work (Schmidgall et al., 2024), while adapting key elements to the context of Arabic medical datasets and healthcare applications. These adaptations include culturally relevant bias categories, prompts designed for clinical scenarios, and additional evaluation metrics to align with real-world needs.
We created a structured framework to systematically assess language-model resistance to cognitive biases:
  1. 1.
    Baseline Testing: Models are evaluated using the original, unbiased dataset to establish baseline performance metrics.
  2. 2.
    Bias Testing: Models are tested with biased variations of prompts. We compute the change in accuracy to assess impact of bias.
  3. 3.
    Evaluation of Bias Mitigation: Mitigation techniques are tested to determine their impact on accuracy. These include:
    • Bias Education: Adding warnings to prompts emphasizing evidence-based reasoning (e.g., “Evaluate each patient uniquely, without relying on trends or recent cases”).
    • One-Shot Demonstration: A single negative example is provided to illustrate incorrect reasoning caused by bias.
    • Few-Shot Demonstration: Both negative and positive examples are presented to show correct and incorrect handling of bias.
This framework provides a systematic and reproducible approach to assess and address cognitive biases in LLMs, ensuring their deployment in healthcare contexts is both effective and ethically sound.
Figure 2: Overall results of SOTA models on the MedArabiQ benchmark. (a) Average performance across benchmark datasets. (b) Breakdown of model performance across benchmark tasks.
Figure 3: Performance Samples from Closed and Open-ended Benchmarks. We report model outputs for samples from (a) Multiple Choice Questions and Fill-in-the-Blank, and (b) Patient-Doctor Q&A tasks, illustrating differences in accuracy, response validity, and language consistency across closed and open-source models.
4 Experimental Setup
Evaluation. We perform zero-shot prompting for all models and tasks, adjusting the temperature according to the nature of each benchmark after an initial testing phase. For closed models, we set the temperature to 0.2 for all tasks. For open-source models, we set the temperature to 0 for Q&A and fill-in-the-blank closed tasks to ensure deterministic answers, and 0.4 for open-ended tasks. For BERTScore evaluation, the model used is XLM-RoBERTa-Large, as it was trained on multiple languages, including Arabic, making it more suitable than monolingual models.
Instruction-tuned Models. We employed instruction-tuned versions of the models described earlier due to their superior ability to interpret and execute task-specific instructions. In contrast, the base versions demonstrated significant limitations in following prompts, even with extensive prompt engineering. This aligns with the literature (Chung et al., 2022; Zhang et al., 2023), which shows that instruction fine-tuning significantly improves model performance and prompt adherence across various tasks and model sizes.
Instruction Prompts. Prompt engineering is crucial for evaluating LLMs. In our prompt engineering experiments, we tested both English and Arabic prompts and found that English prompts were generally more effective. However, for AraMed benchmark tasks, Arabic prompts performed better in open-source models. Based on these findings, we used English prompts for all tasks except patient-doctor Q&A, Q&A with GEC, and Q&A with LLM modifications. Supplementary Table D1 lists the prompts used.
Answer Processing. For Q&A, Q&A with bias, and fill-in-the-blank with choices, models generate both the index of the predicted correct option and the full textual answer. However, some models, particularly open-source ones, exhibit spelling inconsistencies. To ensure accuracy, we evaluated only the first character generated after the phrase “The correct letter is:”, comparing it against valid options. If the character does not match a valid choice, then the response is deemed invalid. For open-ended benchmarks, where answers are not constrained to multiple choice formats, we evaluate the full model-generated response.
Figure 4: Overview of Bias Mitigation Strategy Results for top-performing models. (a) Comparison of accuracy between original questions and questions with bias across different bias categories. The bars represent accuracy percentages, with lighter shades representing a decrease, and darker shades an increase, in accuracy after bias incorporation. (b) Line plot comparing accuracy scores across various mitigation strategies. (c) Radar plot illustrating model performance across different bias mitigation strategies, highlighting relative strengths and weaknesses of each model in handling bias-related challenges.
5 Results
We report all results of the experiments performed on all six benchmarks in Table 1. Our results show that no single model outperforms all others across all benchmarks. For closed tasks (Multiple Choice Questions and fill-in-the-blank with and without choices), closed models perform best, as expected. Gemini 1.5 Pro achieves the highest accuracy in three out of six benchmark tasks, although it performs comparably with Claude 3.5 Sonnet on one. For Multiple Choice Questions and fill-in-the-blank with choices, Gemini performs best, with accuracy scores of 57.5 and 72.7, respectively. Claude 3.5 Sonnet matches Gemini’s 72.7 accuracy on the fill-in-the-blank dataset. In open-ended tasks (Patient-Doctor Q&A, Q&A with GEC, and Q&A with LLM Modifications), Qwen performs best, achieving 85.2 on Patient-Doctor Q&A, while LLaMa scores best on AraMed Q&A with GEC with a score of 85.5. However, models do not consistently excel across both task types, as Qwen and LLaMa perform poorly on closed tasks.
Figure 2 (a) compares the performance of open-source and API-based models averaged across all benchmark tasks. Among all models, Gemini 1.5 Pro achieves the highest average benchmark score, followed closely by GPT-4 and Claude 3.5, reinforcing the dominance of proprietary models in NLP tasks. The open-source models, such as Llama 3.1 and Qwen 2.5, demonstrate competitive but more variable performance, with Qwen 2.5 emerging as the strongest among them. The error bars indicate that while closed models exhibit greater stability across tasks, open-source models show higher performance variability, likely due to their reliance on general-purpose pretraining rather than domain-specific fine-tuning. These findings highlight the current dominance of proprietary models while underscoring the potential of open-source models with further fine-tuning and domain adaptation.
Figure 2 (b) provides a task-specific performance breakdown, revealing significant performance variation across multiple choice, fill-in-the-blank, and generative Q&A tasks. Closed models consistently achieve high scores across all benchmarks, particularly excelling in Q&A tasks, including those with GEC and LLM-enhanced responses. This suggests a strong capability for handling complex medical inquiries, a critical requirement in real-world applications. However, the fill-in-the-blank and multiple choice tasks exhibit more performance divergence, with open-source models like Qwen 2.5 and Llama 3.1 lagging behind the leading closed models. The observed discrepancies emphasize the need for fine-tuning on structured medical datasets, particularly for open-source models, to improve accuracy in knowledge-intensive tasks. Figure 3 shows examples of model responses for the closed and open-ended tasks.
Figure 4 explores the impact of bias and its mitigation strategies on model performance. In Figure 4 (a), we compare the accuracy of GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet-20240620 on original questions versus questions embedded with bias across various bias categories. The results reveal that, generally, all models experience a decline in accuracy when bias is introduced, with the extent of decrease varying by bias type. Gemini 1.5 Pro demonstrates a relatively higher resilience to bias, especially in Status Quo Bias and False-Consensus Bias, while Claude 3.5 Sonnet shows more significant performance drops in Self-Diagnosis Bias and Cultural Bias. Additionally, Figure 4 (b) illustrates the effectiveness of different bias mitigation strategies. Notably, all models exhibit improved accuracy with strategies like Few-Shot prompting compared to biased questions without mitigation. Gemini 1.5 Pro consistently outperforms the other models across these strategies, reinforcing its robustness in handling biased inputs.
Figure 4 (c) presents a radar plot summarizing model performance across the employed bias mitigation strategies. The plot shows Gemini’s consistent performance across all strategies, while GPT-4o and Claude 3.5 Sonnet display larger variability across strategies such as Bias Education and One-Shot prompting. These results further highlight the importance of bias mitigation strategies in enhancing model reliability. Additional information on the performance of models by bias category and medical specialty across different bias mitigation strategies in Figures F1 and G1 in the Appendix.
6 Discussion
Our findings are consistent with prior research on medical LLM benchmarking, where closed-source models outperform open-source models in structured tasks but perform comparably in generative ones. Chen et al. (2025) demonstrated that proprietary models like GPT-4 and Med-PaLM 2 excel in multiple choice and factual retrieval tasks, largely due to stronger pretraining on structured datasets and domain-specific knowledge integration. Similarly, Alonso et al. (2024) found that proprietary models generally achieve higher accuracy in medical question-answering tasks across multiple languages, reinforcing the idea that closed models have better knowledge grounding and factual consistency. Our results support these trends, with Gemini 1.5 Pro and Claude 3.5 Sonnet leading in multiple choice and fill-in-the-blank tasks, suggesting that API-based models are better suited for clinical decision support and structured question-answering tasks.
Figure 4 further illustrates the variability in model performance when exposed to biased content and the effectiveness of different bias mitigation strategies for the top 3 scoring models (GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet-20240620). All models experience an accuracy drop when bias is introduced, with Claude 3.5 Sonnet showing the most pronounced decline across multiple bias categories. Notably, bias mitigation techniques such as few-shot prompting have shown significant improvement in model performance, especially for Gemini 1.5 Pro, which shows the greatest resilience against bias. Yet, no single strategy consistently mitigates bias across all models and categories, underscoring the complexity of bias in medical NLP and the need for further research.
For generative Q&A tasks, however, traditional automatic evaluation metrics such as BERTScore do not fully capture actual model performance. Although GPT-4 and Claude 3.5 produce responses that are highly relevant and contextually accurate as shown in 3, their longer responses result in lower BERTScores compared to ground truth references. This issue has been previously noted by Liu et al. (2025) who found that metrics like ROUGE and BERTScore struggle to effectively assess medical LLMs due to the inherent complexity of medical diagnoses, where multiple treatment options may exist for a single patient. Our findings reinforce this limitation, as models with lower BERTScores sometimes generate high-quality, informative answers that are penalized for verbosity rather than inaccuracy. This discrepancy highlights the need for more nuanced evaluation approaches, including expert human assessment or task-oriented dialogue evaluations, to better reflect real-world usability in medical consultations.
7 Ethical Considerations
Given the sensitive nature of healthcare applications, ethical considerations are paramount when developing and deploying large language models (LLMs). Our benchmark, MedArabiQ, was carefully constructed from publicly available and anonymized educational resources, ensuring compliance with data privacy standards by excluding personally identifiable patient information. Furthermore, we recognize and discuss several ethical implications associated with clinical deployment of LLMs, including potential risks such as misinformation, unintended amplification of biases (cultural, confirmation, or status quo biases), and limited model interpretability. Addressing these concerns, we emphasize the importance of comprehensive validation through hybrid frameworks combining automated evaluation metrics with clinician expert reviews. We advocate for continuous monitoring, clinician oversight, and clearly defined operational guidelines to mitigate potential harm and biases. By explicitly detailing these ethical considerations, we aim to promote responsible and transparent adoption of LLMs, ultimately contributing to safer and more equitable healthcare solutions.”
8 Limitations and Future Work
While our study provides a comprehensive evaluation of Arabic medical LLMs, there are areas that warrant further exploration. First, there is a possibility for contamination in the datasets. Although the past medical exams are not available in structured digital format and required an extensive effort for digitization and cleaning, we cannot completely rule out the possibility of contamination. That being said, model performance can be improved highlighting the importance of the benchmark datasets. As for AraMed, we explicitly incorporated modifications to test for potential memorization considering that it is a publicly available dataset. To verify the validity of the dataset, preliminary results were obtained in collaboration with medical students to assess the Q&A dataset for factuality, relevance, clarity, and complexity on a scale from 1 to 5. The average scores (Relevance: 4.99, Factuality: 4.97, Accuracy: 4.88, Clarity: 4.89) provide initial evidence of the dataset’s reliability and utility. Extensive evaluation is required in future work as the scope and size of the benchmark is expanded.
Additionally, following standard practices in medical NLP, we rely on benchmark datasets rather than live clinical interactions, ensuring reproducibility and ethical compliance. Notably, our Patient-Doctor Q&A dataset is sourced from AraMed (Alasmari et al., 2024), which consists of real consultations conducted on the Altibbi telemedicine platform, ensuring real-world relevance. While live clinical testing could offer additional insights, it presents substantial privacy and regulatory challenges, particularly in the Arab region, where GDPR and HIPAA regulations impose strict limitations on patient data sharing (Theodos and Sittig, 2020). Additionally, many healthcare facilities in the region still rely on paper-based records (Aljawarneh et al., 2024), making large-scale real-time data collection complex. Future work could explore privacy-preserving strategies for integrating real-world clinical assessments in a secure and ethical manner.
Another limitation of our study is the need for more robust bias mitigation techniques. While we evaluate bias susceptibility across multiple tasks, our findings reinforce that existing strategies do not fully eliminate bias, particularly in sensitive clinical decision-making contexts (Alyafeai et al., 2024; Omar et al., 2024). Future research should focus on developing domain-specific bias mitigation approaches that address linguistic and cultural factors unique to Arabic medical NLP.
In this study, we focus on evaluating zero-shot performance, providing an unbiased assessment of how well LLMs perform on Arabic medical tasks without prior adaptation (Kojima et al., 2022). While this establishes a strong baseline, future work could explore fine-tuning on Arabic medical datasets to further enhance domain-specific understanding. Fine-tuning presents an opportunity to improve model accuracy and contextual relevance, though it also requires careful consideration of data availability, computational costs, and generalization risks.
Furthermore, our benchmarks are in Modern Standard Arabic (MSA), following standardization practices to ensure consistency. However, this does not account for dialectal variations, which are commonly used in real-world patient-doctor interactions. Future work could explore incorporating dialectal data to enhance model adaptability across diverse Arabic-speaking healthcare contexts. Additionally, while this study focuses on text-based benchmarks as a unimodal foundation, expanding Arabic medical NLP benchmarks to support multimodal AI, such as medical images and lab results, would be a valuable direction for future research.
In the future, we aim to actively expand MedArabiQ by increasing both the breadth and depth of medical specialties covered. Specifically, we plan to broaden the benchmark to include more specialized clinical domains such as mental health, infectious diseases, and chronic illnesses. This expansion will involve close collaboration with expert physicians and medical educators who will help ensure the questions reflect current clinical standards and practices. Furthermore, we intend to classify questions into specific clinical reasoning types (e.g., diagnostic reasoning, treatment planning, patient counseling), systematically rate their complexity, and conduct rigorous inter-rater reliability assessments. This structured expansion significantly enhances the benchmark’s comprehensiveness, reliability, and utility for fine-tuning Arabic-focused and multilingual LLMs, ultimately improving their clinical applicability and supporting equitable healthcare outcomes for Arabic-speaking populations.
9 Conclusion
In this work, we introduced the first structured benchmark for evaluating LLMs in Arabic healthcare, addressing a critical gap in Arabic medical NLP. Our benchmark consists of 700 diverse clinical samples, covering both structured medical knowledge assessments and real-world patient-doctor interactions. Beyond Arabic healthcare, our benchmark lays the foundation for developing benchmarks in other medically underserved languages, contributing to the global refinement of medical AI applications.
Our evaluation exposes critical limitations in current LLMs, including factual hallucinations in open-ended tasks and vulnerability to biases in clinical decision-making, reinforcing the need for robust bias mitigation strategies. Future work should explore fine-tuning LLMs on Arabic medical datasets, expanding benchmarks to capture dialectal variations, and developing effective bias mitigation strategies tailored to Arabic medical contexts. By releasing our benchmarks, we aim to foster further research in Arabic medical NLP, providing a foundation for trustworthy, unbiased, and effective AI-driven healthcare solutions.
Appendix A Related Work
Several Arabic medical datasets exists, but they each fall short in some way and most are focused on the task of medical question-answering. In Table A1, we summarize these datasets, their limitations, and how we overcome their limitations with our dataset. To summarize, our contribution is four-fold:
  1. 1.
    Introducing a dataset consisting almost entirely of novel proprietary data that has not been seen by language models before, carrying a much lower risk of contamination.
  2. 2.
    Extending data from traditional question answering to more varied use cases through modalities like multiple-choice and fill-in-the-blank, testing medical knowledge and reasoning.
  3. 3.
    Ensuring the quality of data points by manual human curation, inspection, and/or review.
  4. 4.
    Capturing the spirit of real-world patient-doctor interactions using data sourced from the online medical platform Altibbi.
Table A1: Comparison of Various Arabic Medical Question-Answering Datasets
Dataset Name Authors Description Limitations How our paper addresses its limitations
CQA-MD Nakov et al. (2019) A corpus of over 45,000 question-answer pairs sourced from 3 online Arabic medical forums. Though it is comprehensive, more than half of the question-answer pairs were noted as irrelevant to the original question, and almost all are not directly related. This was necessary for the purposes of Nakov et al.’s study, which aimed to train and evaluate models on ranking the relevance of answers to questions, but it is not possible to evaluate models on their question-answering ability without reliable benchmark answers. Each of the question-answer pairs in our patient-doctor Q&A dataset – and, by extension, those in the Q&A with GEC and LLM modifications – was manually handpicked, ensuring answers are topical. The authors of the original dataset, AraMed [3], from which we constructed these three datasets, even performed several preprocessing steps and claim that the dataset “does not contain irrelevant answers.”
ARmed Fehri et al. (2022) A corpus for a proprietary medical question answering (MQA) system. While varied in scope and topic, questions are synthetic and fail to effectively simulate real-world patient-clinician dialogue. Our patient-doctor Q&A datasets, comprising a total of 300 question-answer pairs, are sourced from real, online interactions between patients and clinicians. Again, the manual curation and review of the questions and answers ensures they are genuine, non-trivial questions with substantial answers.
AraMed, Arabic Healthcare Dataset Alasmari et al. (2024), Gawbah (2024) Large-scale Arabic medical question-answering datasets extracted from an online medical forum, Altibbi. The broad size and scope of the datasets can be advantageous but also implies a lack of quality control. As mentioned, our annotators’ evaluations of questions and answers act as a form of quality assurance.
Med42 Saadi et al. (2025) Vast English medical dataset used to train a model, curated from a larger span of resources including chat interactions and chain-of-thought reasoning beyond simple question-answering. The dataset is originally in English but translated to Arabic through an LLM. Saadi et al. themselves found that, with larger models, translations of datasets perform more poorly compared to datasets originally in Arabic. If we were to consider other medical datasets in English, numerous resources exist, but translating them defeats the purpose of our study in regards to introducing original, novel Arabic data. We only use data that is originally in Arabic.
Appendix B Dataset Overview
Table B1 summarizes the datasets that were used to evaluate the models, splitting them by source and describing what each consists of.
[…]
tab:datasets Source Dataset Description AraMed Patient-Doctor Q&A 100 manually selected questions from the AraMed corpus, covering a wide range of medical categories. Q&A with Grammatical Error Correction 100 manually selected questions from the AraMed corpus, with Grammatical Error Correction applied. Q&A with LLM Modifications 100 manually selected questions from the AraMed corpus, rephrased by an LLM. Past Exams Multiple Choice Questions 100 manually selected questions from the past exams. | |and Notes Multiple Choice Questions with Bias 100 manually selected questions from past exams, adapted to test ethical compliance and responses in culturally sensitive scenarios. Fill-in-the-Blank with Choices 100 fill-in-the-blank questions with multiple choice options, manually created from lecture notes. Fill-in-the-Blank without Choices 100 open-ended fill-in-the-blank questions, requiring the generation of answers without prompts.
Table B1: Overview of datasets for evaluating LLMs in Arabic healthcare.
Appendix C Model Overview
Models were categorized in Table C1 based on the transparency of their training data, which affects the possibility of cross-contamination during evaluation, and linguistic coverage, which affects their performance in Arabic.
[…]
tab:model_summary Training Data Transparency Linguistic Coverage Models Unknown Multilingual LLaMA 3.1 70B (Grattafiori et al., 2024) Multilingual Claude 3.5 Sonnet (Anthropic, 2024) Multilingual GPT-4o (OpenAI et al., 2024) Multilingual Qwen 2 72b (Yang et al., 2024a) Multilingual Gemini 1.5 Pro (Team et al., 2024)
Table C1: Summary of selected language models.
Appendix D Prompts by Task
Prompt engineering plays a substantial role in the response produced by the LLM. The same prompt was used for each model, though it was customized to each category of tasks as shown in Table D1. This was crucial to ensure the prompt reflected realistic scenarios as closely as possible, tailored to different use cases.
Ex 6: NR method Ex last: computer first and last derivate and then NR
[…]
tab:task_prompts Benchmark Tasks Prompt Used Multiple Choice Questions, Fill-in-the-Blank (with choices), and Multiple Choice Questions with bias This is a multiple choice question, choose the correct option. The output should consist only of the single letter of the correct answer with no explanation.
{Question:}
{Options:}
Fill-in-the-blank (without choices) Fill in the blank in the question with the correct medical term(s). Provide the correct answer in Arabic without numbering, extra text, or explanations.
{Question:}
{Answer:}
Patient-Doctor Q&A, Q&A with GEC, and Q&A with LLM modifications [Uncaptioned image]
{Question:}
{Answer:}
Table D1: Summary of prompts used for all benchmark tasks.
Appendix E Bias Evaluation
Table E1 compares the performance of models on the multiple choice questions and the multiple choice questions with bias. The performance with bias mitigation strategies, namely bias education, one-shot mitigation, and few-shot mitigation. The accuracy generally decreases with bias and increases again with bias mitigation. More specifically, though bias education decreased accuracy, one-shot mitigation and few-shot mitigation improved performance on all models, with the most substantial increase in accuracy noted for one-shot mitigation on Claude 3.5 Sonnet.
[…]
tab:bias_evaluation_results Model Q&A Q&A with Bias Bias Education One-Shot Mitigation Few-Shot Mitigation Claude 3.5 Sonnet 53.1 36.7 33.6 45.9 41.8 GPT-4o 61.2 43.9 42.9 52 52 Gemini 1.5 Pro 65.3 56.1 54.1 57.1 62.2
Table E1: Comparison of performance by accuracy score without bias, with bias, and with bias mitigation across the 3 Chosen Models.
Appendix F Performance by Bias Category
Figure F1 groups the performance of models by bias category for questions without bias, with bias, and with bias mitigation. Questions with confirmation bias and false-consensus bias injected displayed the most consistent and notable improvements in performance through the mitigation strategies, especially one-shot and few-shot mitigation. Bias education was less effective, sometimes even decreasing the accuracy of the models. Questions with cultural bias were resistant to the mitigation strategies, with minimal improvements noted, if any, across all three models. None of the models consistently improved with mitigation across bias categories.
Figure F1: Model Accuracy Across Bias Categories with Different Mitigation Strategies for Gemini 1.5 Pro, Claude 3.5 Sonnet-20240620, and GPT-4o. (a) Gemini 1.5 Pro. (b) Claude 3.5 Sonnet-20240620. (c) GPT-4o.
Appendix G Performance by Question Category
The relevant field of medicine can be used to group questions while exploring the performance with bias and bias mitigation, as seen in Figure G1. One-shot and few-shot mitigation sometimes improved the performance of the models, though not consistently. Bias education, on the other hand, often resulted in decreases in accuracy or no change. Improvements in accuracy were not conistent at all. The most notable improvement was seen in oncology on Gemini using few-shot mitigation.
Figure G1: Model Accuracy Across Question Categories with Different Mitigation Strategies for Gemini 1.5 Pro, Claude 3.5 Sonnet-20240620, and GPT-4o. (a) Gemini 1.5 Pro. (b) Claude 3.5 Sonnet-20240620. (c) GPT-4o.